OpenStudy (anonymous):

complex numbers chapter : how to find arguments of roots (1/2)+i(square root 3/2) or (1/2) -i(square root 3/2) please help. what is argument in the first place of complex numbers chapter!!

OpenStudy (anonymous):

for christ's sake, someone help !

OpenStudy (anonymous):

(1/2)+i(square root 3/2) ->a = 1/2, b = sqrt 3/2 => tan (theta) = sqrt3 Thus argument is 60 degree ( pi/3)

OpenStudy (anonymous):

@jesonpeson WHERE are you?

OpenStudy (anonymous):

the tan(theta) = sqrt 3 , i didnt understand this part but ur answer is right ! please explain this part ~!!

OpenStudy (anonymous):

Of course, it should be correct :P tan (theta) = b/a

OpenStudy (anonymous):

OHH~!! hahaha thanks alot chlorophyll !! finally. :D cos apparently i'm having my alevel cambridge exams coming this may 16 on P3. thank you.

OpenStudy (anonymous):

My pleasure :)

OpenStudy (anonymous):

muacks~ :p

OpenStudy (anonymous):

It took me a while to figure out the term " root argument" is ANGLE :P